OpenStudy (anonymous):

Find the volume of a cone with radius 30 in and slant height 50 in. A. 3769.91 in3 B. 3768 in3 C. 113040 in3 D. 37699.11 in3

OpenStudy (anonymous):

\[\huge V_{cone}=\frac{1}{3}\pi r h\] the equation for volume of a cone requires height. you can find the height if you consider a slice of the cone as a right angle triangle. 

OpenStudy (anonymous):

Hate to be rude, but I just need the answer...I have 8 minutes and like 7 other hard problems left.

OpenStudy (anonymous):

We tutors can't just give answers. We would get in trouble and also students don't learn anything. Just substitute.

OpenStudy (anonymous):

If I knew how to do it, or was good at it, do you think i'd be asking this...

OpenStudy (anonymous):

4 minutes. Boo. I fail.

OpenStudy (anonymous):

Asking is fine. Then the tutor gives helps and hints, and THEN the student thinks and does some work.

OpenStudy (anonymous):

i think the formula for this one is V=1/3(pi)r^2*h just get the height and substitute

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Awful. No help at all... I've tried before I just get the wrong answer.

OpenStudy (anonymous):

if by awful you mean your reaction to not having answers handed to you, then yes awful indeed

OpenStudy (anonymous):

You've been given lots of help. What is lacking is appreciation of that help.

OpenStudy (anonymous):

Awful, as in your help sucks.

OpenStudy (anonymous):

I'm sure my computer teach could help me better.

OpenStudy (anonymous):

We'll have to have a moderator set you straight on all of this.

OpenStudy (anonymous):

nahhhhhhhhh i run this